PLEASE GIVE STEPS FOR EACH ANSWER! THANK YOU!!!
Answer #1

1. Calculate activity rate :

a. Setup equipment = 270000/5000 = 54 per setup hour

b. other overhead = 900000/6000 = 150 per oven hour

2. Total Overhead assigned to Fudge using ABC :

Total overhead = (4000*54+1000*150) = 366000

3. Unit overhead cost assigned = 366000/5000 = 73.2 per unit

4. Plantwide overhead rate = 1170000/6000 = 195 per oven hour

5.Total overhead assigned to fudge = 195*1000 = 195000